Preparing a Classified Balance Sheet The following balances are from the post-closing trial balance of Armour Inc. Account TiBalance Sheet As of December 31, 2020 Assets Current assets Cash $ 16,000 Accounts receivable $ 65,000 0 Accounts receivable,b. Compute working capital on December 31, 2020. Working capital $ 0 Please answer all parts of the question.
Answer #1

Working capital = Current assets - Current liabilities

Current assets = Cash + account receivable + Inventory + Prepaid insurance + Short term investment - allowance for doubtful account

= 16000 + 65000 + 90000 + 1200 + 28000 - 7500

= 192700

Current liabilities = Account payable + Salaries payable + Interest payable + Income tax payable

= 45000 + 5000 + 1500 + 9000

= 60500

Working capital = Current asset - Current liabilities

= 192700 - 60500 = 132200